OpenStudy (bunnykitten02):

if 5 inches is 100 miles than how many miles is 4 1/20 inches

OpenStudy (bunnykitten02):

@Michele_Laino

OpenStudy (bunnykitten02):

anyone :(

jhonyy9 (jhonyy9):

4 1/20 inches = (4*20+1)/20 = 81/20 use the rule of 3 simples 5 in. ----- 100 miles 81/20 in ----- x miles ------------------- x = (81/20 )*100/5 = (81/20)*20 = (81*20)/20 = ? miles

OpenStudy (ibrafacts):

1 inch equal 20miles

OpenStudy (ibrafacts):

u got 4 inch so 4x 20 =?

OpenStudy (bunnykitten02):

thank you both i would meadal boh if i could

OpenStudy (ibrafacts):

no problem

OpenStudy (ibrafacts):

so whats ur final answer may i ask?

OpenStudy (bunnykitten02):

well i just got kicked out of my lesson darn

OpenStudy (bunnykitten02):

i eep getting different answers 81 84 a few more

OpenStudy (bunnykitten02):

so im not sure